# Arithmetic Properties of Several Generalized-Constant Sequences, with Implications for $\boldsymbol{\Gamma^{\left(n\right)}\left(1\right)}$

### v5

@@ 2.3. Transcendence Results
(ii) We proceed by induction on $m\in\left\{2,3,\ldots,n\right\}$. For the base case, $m=2$, it is known from part (i) that $\mathcal{G}_{2}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\gamma^{\left(4\right)}\right\}$ contains at least one transcendental element. For any $m\geq 2$, we show that if $\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$ contains at least one transcendental element, then it must contain at least two transcendental elements, which in turn implies that $\mathcal{G}_{m+1}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\gamma^{\left(m\right)}\right\}\right)\cup\left\{\gamma^{\left(2m+1\right)},\gamma^{\left(2m+2\right)}\right\}$ contains at least one transcendental element.
@@
(iii) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, part (ii) implies that the set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ contains at least one transcendental element for all $n$. Since $n$ can be arbitrarily large, this precludes the possibility that there exists a finite $n^{*}\in\mathbb{Z}_{\geq 1}$ such that $\gamma^{\left(n\right)}$ is algebraic for all $n\geq n^{*}$.
(iv) As previously observed, if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then part (ii) forces each set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ to contain at least one transcendental element $\gamma^{\left(k_{n}\right)},\>k_{n}\in\left\{n,n+1,\ldots,2n\right\}$. Taking $m=3^{i}$ for $i\in\mathbb{Z}_{\geq 1}$, we now consider the family of disjoint intervals $\left[3^{i},2\cdot 3^{i}\right]\ni k_{m}$ and note that we can choose $N\in\mathbb{Z}_{\geq 2}$ such that $N\geq 2\cdot 3^{i}=2m$ and the $k_{m}\in\left[3^{i},2\cdot 3^{i}\right]\subseteq\left\{1,2,\ldots,N\right\}$ are distinct. This implies that the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N$ is given by
@@
Since Conjecture 1 has been verified computationally for $n\in\left\{2,3,\ldots,26\right\}$, we know from part (ii) of the above theorem that at least two elements of each associated set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental. Furthermore, it is easy to see from the $\textrm{Gumbel}\left(0,1\right)$ moment-generating function, $\textrm{M}_{X}\left(t\right)=\Gamma\left(1-t\right),\>t<1$, that
$\gamma^{\left(n\right)}=\textrm{E}_{X}\left[X^{n}\right]=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right),$
implying that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere. In particular: (a) at least two elements of each set $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(n\right)}\left(1\right),\Gamma^{\left(n+1\right)}\left(1\right),\ldots,\Gamma^{\left(2n\right)}\left(1\right)\right\}$ are transcendental for $n\in\left\{2,3,\ldots,26\right\}$; and (b) if Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, then $\Gamma^{\left(n\right)}\left(1\right)$ must be transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]). [^10] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
@@ 3. Algebraic Independence of the $\eta^{\left(n\right)}$
(I) The $F_{n}\left(z\right)$, like the $w_{n}\left(z\right)$, are $E$-functions (in the sense of Siegel) that satisfy an $\left(n+1\right)$-order non-homogeneous linear differential equation with coefficients in $\mathbb{Q}\left(z\right)$ and unique finite singularity at $z=0$ (see Chapter 7, Section 1 of Shidlovskii [14]).
@@
For the sequence $\left\{\delta^{\left(n\right)}/e\right\}_{n\geq 1}$:
(i) $\delta^{\left(n\right)}/e$ is transcendental infinitely often; and
(ii) the density of the set of transcendental $\delta^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
@@
for all $n$, and begin by noting that this inequality holds for $n<4$ because $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ is generated by at most $|\mathcal{G}_{n}|$ elements, and $|\mathcal{G}_{n}|\leq n\leq n-\lfloor n/2\rfloor+1$ for $n\in\left\{1,2,3\right\}$. In the following analysis, we treat the case of $n\geq 4$.
@@
Given that $\mathcal{B}_{n}$ contains $n-\lfloor n/2\rfloor+1$ elements, the transcendence degree of $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ (which is generated by $\mathcal{B}_{n}$) satisfies (23). Then, from the relation in (8), one can see that $\mathcal{F}_{n}^{\left(\mathcal{H}\right)}$ is contained in the composite field $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right)$, implying
@@ 4.1. Two Additional Sequences
which is shown to be transcendental by Theorem 4 of Section 4.3 below.
@@ 4.3. Transcendence Results
For the sequence $\left\{\widetilde{\delta}^{\left(n\right)}/e\right\}_{n\geq 1}$:
(i) $\widetilde{\delta}^{\left(n\right)}/e$ is transcendental infinitely often; and
(ii) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
@@
The proofs of both parts of this theorem are entirely analogous to those of the corresponding parts of Theorem 3. ∎

### v6

@@ 2.3. Transcendence Results
(ii) For $X\sim\textrm{Gumbel}\left(0,1\right)$, the moment-generating function, $\textrm{M}_{X}\left(t\right)=\textrm{E}_{X}\left[e^{tX}\right],\>t<1$, is given by
$\Gamma\left(1-t\right)={\displaystyle\sum_{n=0}^{\infty}}\textrm{E}_{X}\left[X^{n}\right]\dfrac{t^{n}}{n!}={\displaystyle\sum_{n=0}^{\infty}}\gamma^{\left(n\right)}\dfrac{t^{n}}{n!}.$
Euler’s reflection formula then yields
$\Gamma\left(1+t\right)\Gamma\left(1-t\right)=\dfrac{\pi t}{\sin\left(\pi t\right)}$
$\Longrightarrow\left({\displaystyle\sum_{m=0}^{\infty}}\gamma^{\left(m\right)}\dfrac{\left(-t\right)^{m}}{m!}\right)\left({\displaystyle\sum_{n=0}^{\infty}}\gamma^{\left(n\right)}\dfrac{t^{n}}{n!}\right)={\displaystyle\sum_{k=0}^{\infty}}c_{k}\pi^{2k}t^{2k},$
where $c_{k}\in\mathbb{Q}^{\times}$ for all $k\geq 0$, and we can compare the coefficient of $t^{2k}$ on both sides of the above equation to obtain
${\displaystyle\sum_{j=0}^{2k}}\dfrac{\left(-1\right)^{j}\gamma^{\left(j\right)}\gamma^{\left(2k-j\right)}}{j!\left(2k-j\right)!}=c_{k}\pi^{2k}$ (22)
for all $k$.
Suppose (for purposes of contradiction) that only finitely many $\gamma^{\left(n\right)}$ are transcendental, and choose $n^{*}\in\mathbb{Z}_{\geq 0}$ such that $\gamma^{\left(n\right)}\in\overline{\mathbb{Q}}$ for all $n>n^{*}$, with $\mathcal{W}=\textrm{span}_{\overline{\mathbb{Q}}}\left\{\gamma^{\left(0\right)},\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\ldots,\gamma^{\left(n^{*}\right)}\right\}$ denoting a finite-dimensional $\overline{\mathbb{Q}}$-vector space. Now note that for any $k>n^{*}$, each summand on the left-hand side of (22) has at least one algebraic factor $\gamma^{\left(i_{k}\right)}$, for some $i_{k}>n^{*}$ (because $j+\left(2k-j\right)=2k>2n^{*}\Longrightarrow\left(j>n^{*}\right)\vee\left(2k-j>n^{*}\right)$). It then follows that every summand lies in $\mathcal{W}$, and therefore that $c_{k}\pi^{2k}\in\mathcal{W}$ for all $k>n^{*}$. Since $c_{k}\in\mathbb{Q}^{\times}$, we see that $\pi^{2k}\in\mathcal{W}$ for all $k>n^{*}$, and thus that all $n^{*}+2$ elements $\left\{\pi^{2\left(n^{*}+1\right)},\pi^{2\left(n^{*}+2\right)},\ldots,\pi^{2\left(2n^{*}+2\right)}\right\}$ must lie in the space $\mathcal{W}$. However, the dimension of $\mathcal{W}$ is at most $n^{*}+1$, forcing a non-trivial $\overline{\mathbb{Q}}$-linear dependence among these $n^{*}+2$ elements; that is, a non-trivial polynomial relation over $\overline{\mathbb{Q}}$ satisfied by $\pi^{2}$, which contradicts the transcendence of $\pi$.
(iii) We proceed by induction on $m\in\left\{2,3,\ldots,n\right\}$. For the base case, $m=2$, it is known from part (i) that $\mathcal{G}_{2}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left\{\gamma,\gamma^{\left(2\right)},\gamma^{\left(3\right)},\gamma^{\left(4\right)}\right\}$ contains at least one transcendental element. For any $m\geq 2$, we show that if $\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}$ contains at least one transcendental element, then it must contain at least two transcendental elements, which in turn implies that $\mathcal{G}_{m+1}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}=\left(\left(\mathcal{G}_{m}^{\left(\textrm{U}\right)}\cup\left\{\gamma\right\}\right)\setminus\left\{\gamma^{\left(m\right)}\right\}\right)\cup\left\{\gamma^{\left(2m+1\right)},\gamma^{\left(2m+2\right)}\right\}$ contains at least one transcendental element.
@@
(iv) Assuming Conjecture 1 holds for all $n\in\mathbb{Z}_{\geq 2}$, part (iii) implies that the set $\mathcal{G}_{n}^{\left(\textrm{U}\right)}$ contains at least one transcendental element $\gamma^{\left(k_{n}\right)},\>k_{n}\in\left\{n,n+1,\ldots,2n\right\}$ for all $n$. Taking $m=3^{\ell}$ for $\ell\in\mathbb{Z}_{\geq 1}$, we now consider the family of disjoint intervals $\left[3^{\ell},2\cdot 3^{\ell}\right]\ni k_{m}$ and note that we can choose $N\in\mathbb{Z}_{\geq 2}$ such that $N\geq 2\cdot 3^{\ell}=2m$ and the $k_{m}\in\left[3^{\ell},2\cdot 3^{\ell}\right]\subseteq\left\{1,2,\ldots,N\right\}$ are distinct. This implies that the density of the set of transcendental $\gamma^{\left(n\right)}$ evaluated at $N$ is given by
@@
Since Conjecture 1 has been verified computationally for $n\in\left\{2,3,\ldots,26\right\}$, we know from part (iii) of the above theorem that at least two elements of each associated set $\left\{\gamma,\gamma^{\left(n\right)},\gamma^{\left(n+1\right)},\ldots,\gamma^{\left(2n\right)}\right\}$ are transcendental. Furthermore, the identity $\gamma^{\left(n\right)}=\left(-1\right)^{n}\Gamma^{\left(n\right)}\left(1\right)$ implies that all four parts of Theorem 1 are equally valid if one replaces $\gamma^{\left(\bullet\right)}$ with $\Gamma^{\left(\bullet\right)}\left(1\right)$ everywhere. In particular: (a) at least two elements of each set $\left\{\Gamma^{\left(1\right)}\left(1\right),\Gamma^{\left(n\right)}\left(1\right),\Gamma^{\left(n+1\right)}\left(1\right),\ldots,\Gamma^{\left(2n\right)}\left(1\right)\right\}$ are transcendental for $n\in\left\{2,3,\ldots,26\right\}$; and (b) $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental infinitely often, a result consistent with the common expectation that $\Gamma^{\left(n\right)}\left(1\right)$ is transcendental for all $n\geq 1$ (see, e.g., Rivoal [10] and Fischler and Rivoal [11]). [^10] It may be of interest to some readers to note that the asymptotic lower bound on the density of the set of posited transcendental $\Gamma^{\left(n\right)}\left(1\right)$ is smaller (in magnitude) than the asymptotic lower bound on the density of the set of irrational odd-$n$ $\zeta\left(n\right)$ given by Fischler, Sprang, and Zudilin [12]: $\Omega\left(\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{N/2}\right)=\Omega\left(N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1}\right),$ for arbitrarily small $\varepsilon>0$.
@@ 3. Algebraic Independence of the $\eta^{\left(n\right)}$
(I) The $F_{n}\left(z\right)$, like the $w_{n}\left(z\right)$, are $E$-functions (in the sense of Siegel) that satisfy an $\left(n+1\right)$-order non-homogeneous linear differential equation with coefficients in $\mathbb{Q}\left(z\right)$ and a unique finite singularity at $z=0$ (see Chapter 7, Section 1 of Shidlovskii [14]).
@@
For the sequences $\left\{\delta^{\left(n\right)}/e\right\}_{n\geq 1}$ and $\left\{\delta^{\left(n\right)}\right\}_{n\geq 1}$:
(i) $\delta^{\left(n\right)}/e$ is transcendental infinitely often;
(ii) the density of the set of transcendental $\delta^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$;
(iii) $\delta^{\left(n\right)}$ is transcendental infinitely often; and
(iv) the density of the set of transcendental $\delta^{\left(n\right)}$ (i.e., $\left\{n\leq N:\delta^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
@@
for all $n$, and begin by noting that this inequality holds for $n<4$ because $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ is generated by at most $|\mathcal{G}_{n}|$ elements, and $|\mathcal{G}_{n}|=n\leq n-\lfloor n/2\rfloor+1$ for $n\in\left\{1,2,3\right\}$. In the following analysis, we treat the case of $n\geq 4$.
@@
Given that $\mathcal{B}_{n}$ contains $n-\lfloor n/2\rfloor+1$ elements, the transcendence degree of $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ (which is generated by $\mathcal{B}_{n}$) satisfies (24). Then, from the relation in (8), one can see that
$\mathcal{F}_{n}^{\left(\mathcal{H}\right)}\subseteq\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\mathcal{F}_{n}^{\left(\mathcal{D}\right)}\right),$
implying
@@
(iii) Retain $\mathcal{G}_{n}$ and $\mathcal{F}_{n}^{\left(\mathcal{G}\right)}$ as defined in the proof of part (i), and recall from Theorem 2 that $\eta^{\left(n\right)}=\gamma^{\left(n\right)}+\delta^{\left(n\right)}/e$ is algebraically independent of $\eta^{\left(0\right)}=1-1/e$ for all $n\in\mathbb{Z}_{\geq 1}$. Writing
$\eta^{\left(n\right)}=\gamma^{\left(n\right)}+\delta^{\left(n\right)}\left(1-\eta^{\left(0\right)}\right)$ (26)
as in the proof of Corollary 1(ii), further define $\mathcal{H}_{n}^{\prime}=\left\{\eta^{\left(0\right)},\eta,\eta^{\left(2\right)},\eta^{\left(3\right)},\ldots,\eta^{\left(n\right)}\right\}$ and $\mathcal{D}_{n}^{\prime}=\left\{\delta,\delta^{\left(2\right)},\delta^{\left(3\right)},\ldots,\delta^{\left(n\right)}\right\}$ for $n\in\mathbb{Z}_{\geq 1}$, where $\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}=\mathbb{Q}\left(\mathcal{H}_{n}^{\prime}\right)$ and $\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}=\mathbb{Q}\left(\mathcal{D}_{n}^{\prime}\right)$ denote the corresponding fields over $\mathbb{Q}$. From (26), we know that
$\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}\subseteq\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right),$
which implies
$\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}\right)\leq\mathrm{trdeg}\left(\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right)\right)\leq\mathrm{trdeg}\left(\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\right)+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right).$
Then, since adjoining a single element to a field can increase the transcendence degree by at most one, we have
$\mathrm{trdeg}\left(\mathbb{Q}\left(\eta^{\left(0\right)},\mathcal{G}_{n}\right)\right)=\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\left(\eta^{\left(0\right)}\right)\right)\leq\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{G}\right)}\right)+1\leq n-\lfloor n/2\rfloor+2,$
where the final inequality follows from (24). Moreover, we know from Theorem 2 that the values in $\mathcal{H}_{n}^{\prime}$ are algebraically independent, implying $\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{H}^{\prime}\right)}\right)=|\mathcal{H}_{n}^{\prime}|=n+1$ and therefore
$n+1\leq n-\lfloor n/2\rfloor+2+\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right)$
$\Longleftrightarrow\mathrm{trdeg}\left(\mathcal{F}_{n}^{\left(\mathcal{D}^{\prime}\right)}\right)\geq\lfloor n/2\rfloor-1.$
Since $\underset{n\rightarrow\infty}{\lim}\left(\lfloor n/2\rfloor-1\right)=\infty$, it follows that the sequence $\left\{\delta^{\left(n\right)}\right\}_{n\geq 1}$ must contain an infinite number of transcendental terms.
(iv) The density of the set of transcendental $\delta^{\left(n\right)}$ evaluated at $N\in\mathbb{Z}_{\geq 2}$ satisfies
$\dfrac{\#\left\{n\leq N:\delta^{\left(n\right)}\textrm{ is transcendental}\right\}}{N}\geq\frac{\mathrm{trdeg}\left(\mathcal{F}_{N}^{\left(\mathcal{D}^{\prime}\right)}\right)}{N}\geq\dfrac{\lfloor N/2\rfloor-1}{N},$
with asymptotic lower bound
$\underset{N\rightarrow\infty}{\liminf}\left(\dfrac{\lfloor N/2\rfloor-1}{N}\right)=\dfrac{1}{2}$
as $N\rightarrow\infty$. ∎
@@ 4.1. Two Additional Sequences
which is shown to be transcendental by Theorem 4 of Section 4.3.
@@ 4.2. Asymptotic Behavior
0 $7.3890560989\ldots$ $1.7182818285\ldots$
11 $5.1514643230\ldots$ $1.3179021515\ldots$
22 $11.6100810693\ldots$ $2.2929981451\ldots$
33 $32.2422478187\ldots$ $6.4163856532\ldots$
44 $131.4700021171\ldots$ $24.8036368256\ldots$
55 $651.8492520020\ldots$ $121.9625302861\ldots$
66 $3,916.6763381264\ldots$ $725.7973399920\ldots$
77 $27,400.1009939266\ldots$ $5,060.0849690569\ldots$
88 $219,211.5495238585\ldots$ $40,399.8007638273\ldots$
99 $1,972,830.5386794810\ldots$ $363,237.5069807080\ldots$
$10$ $19,728,269.2785592814\ldots$ $3,630,582.2647192273\ldots$
$11$ $217,010,407.4543390336\ldots$ $39,926,583.2712579089\ldots$
$12$ $2,604,123,546.6077724787\ldots$ $479,060,223.3022788290\ldots$
$13$ $33,853,598,434.1585762781\ldots$ $6,227,401,522.0546076015\ldots$
$14$ $473,950,346,279.9399036234\ldots$ $87,180,954,721.7674533138\ldots$
$15$ $7,109,255,026,416.1913290094\ldots$ $1,307,694,336,767.4617097988\ldots$
@@ 4.3. Transcendence Results
For the sequences $\left\{\widetilde{\delta}^{\left(n\right)}/e\right\}_{n\geq 1}$ and $\left\{\widetilde{\delta}^{\left(n\right)}\right\}_{n\geq 1}$:
(i) $\widetilde{\delta}^{\left(n\right)}/e$ is transcendental infinitely often;
(ii) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}/e$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}/e\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$;
(iii) $\widetilde{\delta}^{\left(n\right)}$ is transcendental infinitely often; and
(iv) the density of the set of transcendental $\widetilde{\delta}^{\left(n\right)}$ (i.e., $\left\{n\leq N:\widetilde{\delta}^{\left(n\right)}\textrm{ is transcendental}\right\}/N$) has asymptotic lower bound $1/2$ as $N\rightarrow\infty$.
@@
The proofs of all parts of this theorem are entirely analogous to those of the corresponding parts of Theorem 3. ∎